through (-1,-1); perpendicular to x+2y=-2
Answer by ewatrrr(24785)

X+ 2Y = -2
Y = (-1/2)X - 1 Green Line
m = -1/2
|New perpendicular Line: m = 2 (negative reciprocal)
thru P(-1,-1)
y + 1 = 2(x+ 1)
y = 2x + 1 Blue Line